# Leveraging Pre-Trained Visual Transformers for Multi-Band Photometric Light Curve Classification

### v2

@@ 4.2 Model performance in ELAsTiCC Challenge
Figure 8: Confusion matrices for the ATAT model trained exclusively on light curves (LC) and the best-performing SwinV2 model, both using the ELAsTiCC dataset. The black lines delineate the three main groups of objects used in the ATAT paper: the Transient, Stochastic, and Periodic groups, from left to right, respectively. Standard deviations are displayed for values greater than or equal to $0.05$.
To assess the model’s performance on a large-scale dataset, we used ELAsTiCC, which includes six bands. In line with prior datasets, we fine-tuned hyperparameters to enhance the model’s effectiveness. The best hyperparameter combination for Grid and Overlay can be found at the bottom of Table 5 in Appendix B.
Table 4 presents the F1-scores along with their standard deviations for SwinV2’s best results, compared to the Astronomical Transformer for time series And Tabular data (ATAT; Cabrera-Vives et al. 2024), a specialized time domain model designed for multi-band light curves. Since ATAT provides results that combine light curves with metadata and additional features, we focused on its light curve-based results to ensure a fair comparison and demonstrate the effectiveness of our approach on equal footing. We will refer to this as ATAT (LC). The results show that SwinV2 slightly outperformed ATAT (LC) in both configurations, with the Overlay approach achieving the highest F1-score of $65.5\%$, followed by the Grid approach with $64.6\%$, while ATAT (LC) obtained $64.2\%$. Although the $1.3\%$ improvement of SwinV2 (Overlay) over ATAT (LC) is moderate, it underscores the model’s ability to effectively distinguish between 20 transient and variable star classes in a more complex classification task.
Figure 9: Distribution of F1-scores obtained during hyperparameter tuning for the ELAsTiCC dataset, comparing the Overlay and Grid multi-band approaches.
@@
Figure 9 illustrates the distribution of F1-scores obtained during the hyperparameter tuning process for the ELAsTiCC dataset, comparing the Overlay and Grid approaches. As before, the median F1-score values across all tested hyperparameter combinations remain consistent with those of the best configurations reported in Table 4, where SwinV2 (Overlay) slightly outperforms SwinV2 (Grid). The F1-score variations across different hyperparameter settings fall within a narrow range of $61\%$ to $67\%$, indicating that classification performance remains relatively stable regardless of the chosen hyperparameters.
@@ 5 Conclusions
We have introduced a simple and effective method for classifying photometric light curves using a pre-trained VT model without the need for complex feature engineering or multi-band preprocessing to handle multi-band observations. We demonstrated the capability and effectiveness of this method using one, two, and six bands of light curves, as well as in scenarios involving imbalanced datasets, limited data, and the classification of both transient and variable stars.
Our results show that leveraging the knowledge acquired from pre-training SwinV2 on a large dataset and fine-tuning it for a specific task can lead to better performance than models specifically designed for light curve data. When using a single band in the MACHO dataset, we demonstrated that, compared to Astromer, our method improves the F1-score by $4.1\%$ in a dataset with 20 spc used to train the model and by $4.4\%$ in a dataset with 500 spc. Additionally, we showed that it is possible to maintain a high performance ($79.8\%$) of F1-score on the Full unbalanced dataset. When incorporating a second band in the same dataset, we observed further improvements compared to SwinV2 with one band: an additional $6.8\%$ in the F1-score for 20 spc, $4.3\%$ for 500 spc, and $4.5\%$ in the Full dataset. Finally, when evaluating our method on the ELAsTiCC dataset, we demonstrated that it achieves an F1-score of $65.5\%$, which is $1.2\%$ higher than the score obtained in ATAT using only light curves.
This study highlights the discriminative power of pre-trained VT models, demonstrating their ability to adapt to datasets that differ entirely from those on which they were originally trained. By analyzing images of light curves, the model successfully distinguished between astronomical objects across datasets with distinct observational characteristics, such as the MACHO survey and the ELAsTiCC challenge. A key factor in this success is the model’s ability to capture the temporal behavior of light curves, which is primarily achieved through the positional embedding. This embedding provides a structured representation of the light curve’s spatial arrangement within the image, allowing the model to implicitly encode ordering information. As a result, it delivers competitive performance against models specifically designed to process sequential light curves and account for the distances between observations. This capability not only underscores the model’s versatility but also suggests a promising direction for developing a more generalizable approach. In particular, it could serve as a foundation for a model capable of adapting to any type of survey with minimal fine-tuning using only a few light curves. However, further investigation is necessary to fully validate this potential.
@@ Appendix A Visual Representations of MACHO and ELAsTiCC Light Curves
Figure 10: Examples from the MACHO dataset illustrating six objects from different classes. Each object is shown in three formats: the original single-band light curve, and two-band representations generated using the Overlay and Grid approaches.
Figure 11: Examples from the ELAsTiCC dataset showing six-band light curves for six objects from different classes. Each object is represented using both the Overlay and Grid approaches.
Figures 10 and 11 present randomly selected examples from the MACHO and ELAsTiCC datasets to illustrate the visual characteristics of the light curves used in our experiments. Each figure shows six objects from different classes, represented using the approaches evaluated in our pipeline and visualized with the hyperparameters that yielded the best classification performance. For MACHO, we include three visual representations per object: the single-band light curve (One-band), and the two-band inputs constructed using the Overlay and Grid approaches. For ELAsTiCC, which contains six-band light curves, we show the corresponding Overlay and Grid representations. Within each dataset, the same objects and class labels are used across approaches to enable direct visual comparison of how different methods encode temporal and spectral information.
@@ Appendix B Hyperparameters
Table 5 presents the best hyperparameters obtained for marker size, line width, flux errors, input format, and learning rate for the MACHO dataset (single-band and two-band) in the scenarios of 20 spc, 500 spc, and the Full dataset, as well as for the ELASTiCC dataset (six-band). Hyperparameter tuning was conducted using a grid search over the values listed in Table 1.
The results reveal a few key patterns: the optimal marker size and line width depend on both the number of bands and the approach used to represent multi-band observations. When using the Grid approach, a larger marker size is necessary because each light curve band is represented on a lower-resolution grid, requiring larger observation markers for better visibility. This effect is particularly evident in ELAsTiCC, where the grid is divided into six sections. The same pattern is observed in the Full dataset of MACHO (two-band) but not in the 20 spc and 500 spc scenarios, where the limited number of observations increases the likelihood of weak data representation. Conversely, the Overlay approach generally exhibits less variation in the marker size and line width hyperparameters across the different Full datasets, including MACHO single-band, MACHO two-band, and ELAsTiCC, remaining within a marker size range of 1.0 to 2.0 and a line width range of 0.5 to 2.0. This consistency occurs because the resolution at which bands are represented remains fixed at $256\times 256$. In the same context, flux errors appear to be important in almost all cases, except in the 500 spc scenario for MACHO (two-band). Additionally, higher learning rates performed better for smaller datasets, while lower learning rates were more effective for larger datasets. This trend likely occurs because higher learning rates help escape local optima when data is scarce, whereas lower learning rates enable more stable updates when training on larger datasets.

### v4

@@ 4.2 Model performance in ELAsTiCC
Figure 8: Confusion matrices for the ATAT model trained exclusively on light curves and the best-performing SwinV2 model, both of which use the ELAsTiCC dataset. The black lines delineate the three main groups of objects used in the ATAT paper: the transient, stochastic, and periodic groups, from left to right, respectively. Standard deviations are displayed for values greater than or equal to $0.05$.
To assess the model’s performance on a large-scale dataset, we used ELAsTiCC, which includes six bands. In line with prior datasets, we fine-tuned hyperparameters to enhance the model’s effectiveness. The best hyperparameter combination for “Grid” and “Overlay” can be found at the bottom of Table 5 in Appendix B.
Table 4 presents the F1-scores along with their standard deviations for SwinV2’s best results, compared to the Astronomical Transformer for time series And Tabular data (ATAT; Cabrera-Vives et al. 2024), a specialized time domain model designed for multi-band light curves. Since ATAT provides results that combine light curves with metadata and additional features, we focused on its light curve-based results to ensure a fair comparison and demonstrate the effectiveness of our approach on equal footing. We refer to this as ATAT (LC). The results show that SwinV2 slightly outperformed ATAT (LC) in both configurations, with the “Overlay” approach achieving the highest F1-score of $65.5\%$, followed by the “Grid” approach with $64.6\%$, while ATAT (LC) obtained $64.2\%$. Although the $1.3\%$ improvement of SwinV2 (Overlay) over ATAT (LC) is moderate, it underscores the model’s ability to effectively distinguish between 20 transient and variable star classes in a more complex classification task.
Figure 9: Distribution of F1-scores obtained during hyperparameter tuning for the ELAsTiCC dataset comparing the “Overlay” and “Grid” multi-band approaches.
@@
Figure 9 illustrates the distribution of F1-scores obtained during the hyperparameter tuning process for the ELAsTiCC dataset, comparing the “Overlay” and “Grid” approaches. As before, the median F1-score values across all tested hyperparameter combinations remain consistent with those of the best configurations reported in Table 4, where SwinV2 (Overlay) slightly outperforms SwinV2 (Grid). The F1-score variations across different hyperparameter settings fall within a narrow range of $61\%$ to $67\%$, indicating that classification performance remains relatively stable regardless of the chosen hyperparameters.
@@ 5 Conclusions
We have introduced a simple and effective method for classifying photometric light curves using a pre-trained VT model without the need for complex feature engineering or multi-band preprocessing to handle multi-band observations. We demonstrated the capability and effectiveness of this method using one, two, and six bands of light curves as well as scenarios involving imbalanced datasets, limited data, and the classification of both transient and variable stars.
Our results show that leveraging the knowledge acquired from pre-training SwinV2 on a large dataset and fine-tuning it for a specific task can lead to better performance than models specifically designed for light curve data. When using a single band in the MACHO dataset, we demonstrated that, compared to Astromer, our method improves the F1-score by $4.1\%$ in a dataset with 20 spc used to train the model and by $4.4\%$ in a dataset with 500 spc. Additionally, we showed that it is possible to maintain a high performance ($79.8\%$) of F1-score on the unbalanced “full dataset”. When incorporating a second band in the same dataset, we observed further improvements compared to SwinV2 with one band: an additional $6.8\%$ in the F1-score for 20 spc, $4.3\%$ for 500 spc, and $4.5\%$ in the “full dataset”. Finally, when evaluating our method on the ELAsTiCC dataset, we demonstrated that it achieves an F1-score of $65.5\%$, which is $1.2\%$ higher than the score obtained in ATAT using only light curves.
This study highlights the discriminative power of pre-trained VT models and demonstrates their ability to adapt to datasets that differ entirely from those on which they were originally trained. By analyzing images of light curves, the model successfully distinguished between astronomical objects across datasets with distinct observational characteristics, such as the MACHO survey and ELAsTiCC. A key factor in this success is the model’s ability to capture the temporal behavior of light curves, which is primarily achieved through the positional embedding. This embedding provides a structured representation of the light curve’s spatial arrangement within the image, allowing the model to implicitly encode ordering information. As a result, it delivers competitive performance against models specifically designed to process sequential light curves and account for the distances between observations. This capability not only underscores the model’s versatility but also suggests a promising direction for developing a more generalizable approach. In particular, it could serve as a foundation for a model capable of adapting to any type of survey with minimal fine-tuning using only a few light curves. However, further investigation is necessary to fully validate this potential.
@@ Appendix A Visual representations of MACHO and ELAsTiCC light curves
Figure 10: Examples from the MACHO dataset illustrating six objects from different classes. Each object is shown in three formats: the original single-band light curve and two-band representations generated using the “Overlay” and “Grid” approaches.
Figure 11: Examples from the ELAsTiCC dataset showing six-band light curves for six objects from different classes. Each object is represented using both the “Overlay” and “Grid” approaches.
Figures 10 and 11 present randomly selected examples from the MACHO and ELAsTiCC datasets to illustrate the visual characteristics of the light curves used in our experiments. Each figure shows six objects from different classes, represented using the approaches evaluated in our pipeline and visualized with the hyperparameters that yielded the best classification performance. For MACHO, we include three visual representations per object: the single-band light curve (One-band), and the two-band inputs constructed using the “Overlay” and “Grid” approaches. For ELAsTiCC, which contains six-band light curves, we show the corresponding “Overlay” and “Grid” representations. Within each dataset, the same objects and class labels are used across approaches to enable direct visual comparison of how different methods encode temporal and spectral information.
@@ Appendix B Hyperparameters
Table 5 presents the best hyperparameters obtained for marker size, line width, flux errors, input format, and learning rate for the MACHO dataset (single-band and two-band) in the scenarios of 20 spc, 500 spc, and the “full dataset”, as well as for the ELASTiCC dataset (six-band). Hyperparameter tuning was conducted using a grid search over the values listed in Table 1.
The results reveal a few key patterns: the optimal marker size and line width depend on both the number of bands and the approach used to represent multi-band observations. When using the “Grid” approach, a larger marker size is necessary because each light curve band is represented on a lower-resolution grid, requiring larger observation markers for better visibility. This effect is particularly evident in ELAsTiCC, where the grid is divided into six sections. The same pattern is observed in the “full dataset” of MACHO (two-band) but not in the 20 spc and 500 spc scenarios, where the limited number of observations increases the likelihood of weak data representation. Conversely, the “Overlay” approach generally exhibits less variation in the marker size and line width hyperparameters across the different “full datasets”, including MACHO single-band, MACHO two-band, and ELAsTiCC, remaining within a marker size range of 1.0 to 2.0 and a line width range of 0.5 to 2.0. This consistency occurs because the resolution at which bands are represented remains fixed at $256\times 256$. In the same context, flux errors appear to be important in almost all cases, except in the 500 spc scenario for MACHO (two-band). Additionally, higher learning rates performed better for smaller datasets, while lower learning rates were more effective for larger datasets. This trend likely occurs because higher learning rates help escape local optima when data is scarce, whereas lower learning rates enable more stable updates when training on larger datasets.